OpenStudy (anonymous):

If F(x)=x2⋅ex, what is the exact value of F′(−2)?

OpenStudy (anonymous):

is it e.x or e raise to the power of x????

OpenStudy (anonymous):

x^2 and e^x

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

my pleasure

OpenStudy (beccaboo333):

Hi, and welcome to OpenStudy! I see you're getting help already. But remember this is a Mathematics question. You asked it in English. Please ask next time in the correct section. Thank you!

OpenStudy (anonymous):

lol sorry first time on this thing. will do

OpenStudy (beccaboo333):

If you have any any questions about this site you can pm me.

OpenStudy (anonymous):

alright thanks!

OpenStudy (anonymous):

Yes, see math section. F(x) = (x^2) exp(x) F'(x) = (2x) exp(x) + (x^2) exp(x) F'(-2) = (-4) exp(-2) + (4) exp(-2) = 0 So answer obtained above was correct.